379 views
7 7 votes

A multiple choice exam has $4$ choices for each question. The student has studied enough so that the probability they will know the answer to a question is $0.5$, the probability that the student will be able to eliminate one choice is $0.25$, otherwise all $4$ choices seem equally plausible. If they know the answer they will get the question correct. If not, they have to guess from the $3$ or $4$ choices. As the teacher you would like the test to measure what the student knows, and not how well they can guess. If the student answers a question correctly, what is the probability that they actually know the answer?

  1. $\dfrac{24}{31}$
     
  2. $\dfrac{1}{2}$
     
  3. $\dfrac{31}{48}$
     
  4. $\dfrac{12}{31}$

3 Answers

0 0 votes

Let $K$ be the event that the student knows the answer.  
Let $E$ be the event that the student can eliminate one choice.  
Let $G$ be the event that the student guesses from all 4 choices.  
Let $C$ be the event that the student answers correctly.  

From the problem description, we have the following probabilities:

  • $P(K)=0.5=\dfrac{1}{2}$
     
  • $P(E)=0.25=\dfrac{1}{4}$
     
  • $P(G)=1-P(K)-P(E)=1-0.5-0.25=0.25=\dfrac{1}{4}$
     

The conditional probabilities of getting the correct answer given their state of knowledge are:

  • $P(C\mid K)=1$ (If they know the answer, they get it right)
     
  • $P(C\mid E)=\dfrac{1}{3}$ (Guessing from the $3$ remaining choices)
     
  • $P(C\mid G)=\dfrac{1}{4}$ (Guessing from all $4$ choices)
     

We want to find $P(K\mid C)$, the probability that the student knows the answer given that they answered correctly. We use Bayes’ Theorem:

\[
P(K\mid C)=\frac{P(C\mid K)P(K)}{P(C)}
\]

First, calculate the total probability of a correct answer, $P(C)$, using the Law of Total Probability:

\[
P(C)=P(C\mid K)P(K)+P(C\mid E)P(E)+P(C\mid G)P(G)
\]

\[
P(C)=\left(1\cdot \frac{1}{2}\right)+\left(\frac{1}{3}\cdot \frac{1}{4}\right)+\left(\frac{1}{4}\cdot \frac{1}{4}\right)
\]

\[
P(C)=\frac{1}{2}+\frac{1}{12}+\frac{1}{16}
\]

To add these fractions, find the least common multiple of $2, 12,$ and $16,$ which is $48:$

\[
P(C)=\frac{24}{48}+\frac{4}{48}+\frac{3}{48}=\frac{31}{48}
\]

Now, substitute this back into our Bayes’ Theorem equation:

\[
P(K\mid C)=\frac{1\cdot \frac{1}{2}}{\frac{31}{48}}
=\frac{\frac{1}{2}}{\frac{31}{48}}
=\frac{1}{2}\cdot \frac{48}{31}
=\frac{24}{31}
\]

The correct option is A.

Answer:
Position:
Show:

Related questions

5 5 votes
4 4 answers
390
390 views
GO Classes asked Apr 7
390 views
There are three doors; behind one is a nice car, and behind each of the other two is a goat eating a bale of straw. You choose a door. Then Monty Hall opens one of the ot...
3 3 votes
2 2 answers
261
261 views
GO Classes asked Apr 7
261 views
Sarah and Bob draw $13$ cards each from a standard deck of $52$ cards. Given that Sarah has exactly two aces, what is the probability that Bob has exactly one ace?$\dfrac...
3 3 votes
3 3 answers
275
275 views
GO Classes asked Apr 7
275 views
Consider a clinical test for cancer that can yield either a positive $(+)$ or negative $(-)$ result. Suppose that a patient who truly has cancer has a $1\%$ chance of sli...
3 3 votes
2 2 answers
273
273 views
GO Classes asked Apr 7
273 views
Mark all expressions which indicate that $\mathrm{X}$ is independent of $\mathrm{Y}$ given $\mathrm{Z}$.$P(X, Y \mid Z)=P(X \mid Z) P(Y \mid Z) $ $P(X, Y, Z)=P(X, Z) P(Y)...